OpenStudy (anonymous):

limit x approaches infinity, (rad(x^2-9))/(2x-6)

OpenStudy (anonymous):

come back @hartnn

hartnn (hartnn):

lol! whats rad ?

OpenStudy (anonymous):

sqrt

OpenStudy (anonymous):

ya

hartnn (hartnn):

divide x in numerator and denom.

OpenStudy (anonymous):

5x-6

OpenStudy (anonymous):

rad(x^2-9) ---------- ----> multiply by this rad(x^2-9)

OpenStudy (anonymous):

-y

OpenStudy (anonymous):

then what algebraic

OpenStudy (anonymous):

r2+s2-t2

OpenStudy (anonymous):

x^2-9 : difference of squares what are the factors?

hartnn (hartnn):

@Algebraic! is there any need to do that ?? x-> infinity and not 3.

OpenStudy (anonymous):

ya whoops.

OpenStudy (anonymous):

easy mode

OpenStudy (anonymous):

so im here now (x-3)(x+3)/(2x-6)(sqrt(x^2-9))

OpenStudy (anonymous):

don't bother with what I said.... for limit x-> infinity it's just sqrt(x^2) / 2x

OpenStudy (anonymous):

ok so starting over, i factor whats in the sqrt then factor the bottom, then cancel out whatever i can \[\frac{ \sqrt{(x-3)(x+3)} }{ 2(x-3) }\] \[\frac{ \sqrt{(x+3)} }{ 2 }\]

OpenStudy (anonymous):

hmm, that's not allowed... but here's an easy way to think about it

OpenStudy (anonymous):

imagine putting in x=one million or so...

OpenStudy (anonymous):

the -9 becomes insignificant the -6 becomes insignificant you're left with rad( 1,000,000^2) (which is just 1,000,000) divided by 2,000,000

OpenStudy (anonymous):

so i can pick any number basically and it will be 1/2

OpenStudy (anonymous):

any large number aka x-> infinity

OpenStudy (anonymous):

algebraically:

OpenStudy (anonymous):

\[\frac{ \sqrt{x^{2}(1-\frac{ 9} { x ^{2} })} }{ x(2-\frac{ 6 }{ x} )}\]

hartnn (hartnn):

^^ thats the correct way to solve.

OpenStudy (anonymous):

factor out x^2 from the term under the radical factor out x from the term in the denom (as hartnn was saying before I cut him off)

OpenStudy (anonymous):

\[\frac{ x \sqrt{(1-\frac{ 9} { x ^{2} })} }{ x(2-\frac{ 6 }{ x}) }\]

OpenStudy (anonymous):

x/x cancels as x gets large 9/x^2 -> 0 6/x -> 0 all that's left is rad(1) /2 = 1/2

OpenStudy (anonymous):

thats an odd way but makes sense lol thanks